OpenStudy (anonymous):

Marcus painted 1/3 of his bedroom in 4/5 of an hour. At this rate, how long would it take him the paint the entire room? 5/12 of an hour 12/5 of an hour 15/4 of an hour 4/15 of an hour

OpenStudy (anonymous):

@Atsie

OpenStudy (anonymous):

123456789. Glad to help :)

OpenStudy (anonymous):

@jabez177

OpenStudy (retireed):

Wow I was way off on this answer. It is quite simple. if he paints 1/3 in 4/5 hour, he will paint another 1/3 in another 4/5 hour and he will paint the last third in another 4/5 hour. so all combined 4/5 + 4/5 + 4/5 would be the total time or three times 4/5 hours

jabez177 (jabez177):

\[\frac{ 3 }{ 1 }\times \frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

thanks